OpenStudy (anonymous):

A line has slope 2 and (3, 4) is on the line. What's the correct equation that describes this line in slope-intercept form?

OpenStudy (mathstudent55):

Slope-intercept form: y = mx + b m = slope, b = y-intercept You have the slope, m = 2, and you know one point, (3, 4). Write as much of the equation as you know already: y = 2x + b You still don't know b, but you can find it. Replace x and y in y = 2x + b with the point you are given and solve for b. Then rewrite the equation weith the given slope and the value you find for b.

OpenStudy (anonymous):

So it'd be y=2x - 2 ?

OpenStudy (mathstudent55):

You have y = 2x + b Substitute x and y with 3 and 4, repectively. What equation do you get?

OpenStudy (anonymous):

It's either 2x=y +2 or y=x + 2

OpenStudy (mathstudent55):

I'm not interested in choices. I'm interested in learning how to solve the problem. We were up to y = 2x + b Now replace x with 3 and y with 4. What equation do you get?

OpenStudy (anonymous):

do you multiply 2 and 3 ? if your replacing the x with three ?

OpenStudy (mathstudent55):

Can you please just replace x with 3 and y with 4 and show it?